# RegD: Hierarchical Embeddings via Distances over Geometric Regions

### v1

@@ 3.2 Comparasion with the Hyperbolic Distance
Therefore, by using only elementary arithmetic operations, our depth distance shares similar embedding capabilities to those of hyperbolic spaces. This offers two key advantages over hyperbolic approaches: (1) our method eliminates the need for high-precision numerical representations, such as double-precision tensors (DoubleTensor), which can increase computational complexity and memory usage; and (2) our approach does not require specialized training strategies tailored to the hyperbolic space, which are often essential for achieving optimal performance in hyperbolic space models.
@@
Figure 2: Illustration of the boundary distances between $\textit{reg}_{1}$ and $\textit{reg}_{2}$ in two cases: $\textit{reg}_{2}\subseteq\textit{reg}_{1}$ (left) and $\textit{reg}_{2}\not\subseteq\textit{reg}_{1}$ (right).
4 Boundary Distance: Non-Symmetric Distance for Inclusion
Although the depth distance $d_{\text{dep}}$ introduced above has been shown to have great power for embedding hierarchical data, it is a symmetric distance and therefore inadequate for fully capturing the inherently non-symmetric hierarchical relationships between objects. To address this limitation, we introduce the boundary distance, specifically designed to reflect the partial order of regions defined by set inclusion.
To the best of our knowledge, the concept of boundary distance was first introduced and applied to the embedding of hierarchical data by ShadowCone (Yu et al., 2024). However, their construction is limited to specific types of cones defined in hyperbolic space, and its computation requires specialized functions (e.g., $\operatorname{arcsinh}$), which introduce additional computational costs. In contrast, we propose a more general construction of the boundary distance that can be applied to arbitrary regions in the Euclidean space. Furthermore, our approach is computationally efficient for commonly used geometric objects, such as balls and boxes, requiring only basic arithmetic operations.
4.1 Construction
The boundary distance is defined to measure the minimal cost associated with transforming the spatial relationship between two regions $\textit{reg}_{1}$ and $\textit{reg}_{2}$. Specifically, it quantifies the cost of moving $\textit{reg}_{2}$ out of $\textit{reg}_{1}$ when $\textit{reg}_{2}\subseteq\textit{reg}_{1}$, or moving $\textit{reg}_{2}$ into $\textit{reg}_{1}$ otherwise (when $\textit{reg}_{2}\not\subseteq\textit{reg}_{1}$). This cost can be defined for arbitrary geometric objects based on either distance or volume within Euclidean or other spaces. Below, we introduce a boundary distance based on the Euclidean distance for two regions $\textit{reg}_{1},\textit{reg}_{2}\subseteq\mathbb{R}^{n}$, which consists of two cases:
  1. 1.
    Containment ($\textit{reg}_{2}\subseteq\textit{reg}_{1}$): As illustrated in the left of Figure 2, when $\textit{reg}_{2}$ is fully contained within $\textit{reg}_{1}$, the boundary distance is defined by the minimum Euclidean distance between the complementary region $\textit{reg}_{1}^{c}$ and the points in $\textit{reg}_{2}$ (i.e., length of the red line). This distance quantifies the minimum translation cost required to move at least a part of $\textit{reg}_{2}$ out of $\textit{reg}_{1}$.
  2. 2.
    Non-Containment ($\textit{reg}_{2}\not\subseteq\textit{reg}_{1}$): As shown in the right of Figure 2, when $\textit{reg}_{2}$ is not fully contained within $\textit{reg}_{1}$, the boundary distance is defined as the maximum Euclidean distance from the points in $\textit{reg}_{2}\setminus\textit{reg}_{1}$ to $\textit{reg}_{1}$ (i.e., length of the red line). This distance quantifies the minimum translation cost for moving $\textit{reg}_{2}$ into $\textit{reg}_{1}$.
Let $d(\mathbf{x},\textit{reg}):=\min\{||\mathbf{x}-\mathbf{y}||_{2}\mid\mathbf{y}\in\textit{reg}\}$ be the distance of a points $\mathbf{x}$ to a region reg defined by the minimal distance from $\mathbf{x}$ to $\mathbf{y}\in\textit{reg}$. The formal definition of the boundary distance is as follows:
Definition 2 (Boundary Distance).
Given a region space $\mathcal{R}$, we define the boundary distance over $\textit{reg}_{1},\textit{reg}_{2}\in\mathcal{R}$ by:
@@ 4.1 Construction
Note that a negative sign is added to $d_{\text{bd}}(\textit{reg}_{1},\textit{reg}_{2})$ in the containment case ($\textit{reg}_{2}\subseteq\textit{reg}_{1}$) to clearly distinguish it from other cases. Moreover, the boundary distance is inherently asymmetric, that is, $d_{\text{bd}}(\textit{reg}_{1},\textit{reg}_{2})\neq d_{\text{bd}}(\textit{reg}_{2},\textit{reg}_{1})$ in general. For example, as illustrated in Figure 2, the boundary distance in the reverse order, $d_{\text{bd}}(\textit{reg}_{2},\textit{reg}_{1})$, corresponds to the length of the green dashed line, which differs from the red line representing $d_{\text{bd}}(\textit{reg}_{1},\textit{reg}_{2})$.
Moreover, for widely used geometric objects such as balls and boxes, the boundary distance can be computed efficiently using simple arithmetic operations.
@@
If $\textit{reg}_{1}=\textit{ball}(\mathbf{c}_{1},r_{1})$ and $\textit{reg}_{2}=\textit{ball}(\mathbf{c}_{2},r_{2})$, the boundary distance have the form:
@@
For the case of boxes, where $\textit{reg}_{1}=\textit{box}(\mathbf{c}_{1},\mathbf{o}_{1})$ and $\textit{reg}_{2}=\textit{box}(\mathbf{c}_{2},\mathbf{o}_{2})$, the boundary distance have the form:
$d_{\text{bd}}($ $\textit{reg}_{1},\textit{reg}_{2})=$
$\begin{cases}\max(|\mathbf{c}_{1}-\mathbf{c}_{2}|+\mathbf{o}_{2}-\mathbf{o}_{1})&\textit{ if }\textit{reg}_{2}\subseteq\textit{reg}_{1},\\ ||\max\{|\mathbf{c}_{1}-\mathbf{c}_{2}|+\mathbf{o}_{2}-\mathbf{o}_{1},\mathbf{0}\}||_{2}&\textit{else}.\end{cases}$
@@
The following proposition demonstrates that our definition of the boundary distance effectively captures the inclusion relationship between two regions in two key aspects, as illustrated in Figure 3: (i) it identifies whether one region is (exactly) contained within another, and (ii) it enhances discrimination in inclusion chains, as smaller regions tend to have larger boundary distances. This property is useful for distinguishing shallow children from deeper ones. It is worth noting that the proposition below applies to any regions, not only boxes or balls.
@@
For the boundary distance $d_{\text{bd}}$ defined in Definition 2, the following properties hold:
@@
    If $\textit{reg}_{2}^{\prime}\subseteq\textit{reg}_{2}\subseteq\textit{reg}_{1}$, then
$\textit{reg}_{2}$$\textit{reg}_{1}$$\textit{reg}_{1}$$\textit{reg}_{2}$$\textit{reg}_{2}^{\prime}$
4.2 Specific Constructions for Boxes or Balls
For specific geometric regions like balls and boxes, we can create specialized distance functions to measure the set-inclusion relationship based on their intrinsic geometric properties or established methods. Our framework accommodates these specialized metrics by allowing them to replace the general boundary distance function.
@@ 4.2 Specific Constructions for Boxes or Balls
    Volume-based distance for boxes: Since the volume of a box can be computed as the product of its offsets along different dimensions, we can define a partial distance based on volume:
@@
    The negative logarithm ensures that the distance is non-negative and equals zero only when $\textit{reg}_{2}\subseteq\textit{reg}_{1}$.
@@
    hyperbolic distance for balls: Yu et al. (2024) introduced a series of circular cones in hyperbolic space and defined a boundary distance based on the hyperbolic distance between the apex of these cones. By utilizing the natural mapping from balls to circular cones, we can derive a new boundary distance for balls as follows:
    $d_{\text{bd}}^{\text{cone}}(\textit{reg}_{1},\textit{reg}_{2})=\operatorname{arcsinh}\left(\frac{||\mathbf{c}_{1}-\mathbf{c}_{2}||_{2}-r_{1}}{r_{2}}\right)+a,$ (6)
    where $a=\operatorname{arcsinh}(1)$. See Appendix B for more details.
5 Training Strategy
Energy
For a given pair $(u,v)$, we define their energy as:
@@ 5 Training Strategy
where $\lambda$ is a weighting parameter that balances the contributions of the hyperbolic-like depth distance. This formulation ensures the simultaneous preservation of hyperbolic structural properties when needed, while maintaining the asymmetry characteristic of hierarchical relationships.
Loss
To train our model, we employ a modified contrastive loss function as in (Yu et al., 2024):
$\mathcal{L}(\gamma_{1},\gamma_{2})=$ $\sum_{(u,v)\in P}\Bigg{(}\max\{E(u,v),\gamma_{1}\}$ (8)
$+\log\bigg{(}\sum_{(u,v^{\prime})\in N}e^{\max\{\gamma_{2}-d_{\text{bd}}(\textit{reg}_{u},\textit{reg}_{v^{\prime}}),0\}}\bigg{)}\Bigg{)},$
where $P$ and $N$ denote positive and negative sample pairs, respectively. For positive pairs $(u,v)\in P$, the loss based on $E(u,v)$ promotes both the containment of $\textit{reg}_{v}$ within $\textit{reg}_{u}$ (via the $d_{\text{bd}}$ term) and their geometric similarity (via the $d_{\text{dep}}$ term), whose contributions are controlled by the weight $\lambda$ and the threshold $\gamma_{1}$. For negative pairs $(u,v^{\prime})\in N$, since our primary goal is to push $\textit{reg}_{v^{\prime}}$ outside of $\textit{reg}_{u}$, it is sufficient to use the boundary distance $d_{\text{bd}}(\textit{reg}_{u},\textit{reg}_{v^{\prime}})$ rather than the energy $E(u,v^{\prime})$. A threshold $\gamma_{2}$ is used to regulate how far $\textit{reg}_{v^{\prime}}$ is pushed from $\textit{reg}_{u}$.
Finally, we say that $u$ is considered a parent of $v$ (i.e., $u\prec v$) if $E(u,v)\leq t$, where $t$ is a threshold that achieved the best performance on the evaluation set.
@@
ShadowCone (Umbral-half-space) 57.7 69.4 45.2 52.2 36.8 40.1 32.8 32.6
(Penumbralhalf-space) 52.8 67.8 44.6 51.7 35.0 37.6 26.8 28.4
RegD (box) 64.9 74.4 53.8 51.3 50.7 58.5 42.8 49.6
@@ 6.1 Inferences over DAG
Following (Yu et al., 2024), we evaluate our method on four real-world datasets consisting of Is-A relations: MCG (Wang et al., 2015; Wu et al., 2012), Hearst patterns (Hearst, 1992), the WordNet (Fellbaum, 1998) noun taxonomy, and its mammal subgraph. All models are trained exclusively on basic edges, which are edges not implied transitively by other edges in the graph. For validation and testing, we use the same sets as in (Yu et al., 2024), consisting of 5% of non-basic (inferred) edges, ensuring a fair comparison.
We exclude non-basic edges from training since they can be transitively derived from basic edges. Including them would artificially inflate performance metrics without properly evaluating the embeddings’ ability to capture hierarchical structures. For completeness, results for non-basic cases are provided in Appendix F.1.
Table 2: Summary of Graph Metrics, where $\delta$-Hyp (max/mean) indicates the $\delta$-Hyperbolicity (see Appendix D). Lower values suggest more “hyperbolic” structures, and a value of 0 indicates a tree.
Mammal Noun MCG Hearst
Nodes 1,179 82,114 22,665 35,545
Edges 1,176 84,363 38,288 42,423
$\delta$-Hyp 0.0 / 0.0 4.5 / 0.512 2.0 / 0.316 2.5 / 0.426
Baselines
@@
We compare our method RegD with (i) hyperbolic approaches such as EntailmentCone (Ganea et al., 2018) and ShadowCone (Yu et al., 2024), which is the latest method with the state-of-the-art performance; (ii) region-based methods like OrderEmbedding (Bordes et al., 2013), and tBox (Boratko et al., 2021). We also compare with the ontology embedding methods, ELBE (Peng et al., 2022) and ELEM (Kulmanov et al., 2019), which can be considered as the baseline approaches embedding the DAG as boxes or balls, respectively. As in previous studies, the performance is evaluated using F1-scores.
Results
@@
The performance comparison across different DAGs is shown in Table 1. RegD achieved the best performance on all four datasets. Notably, the box variant consistently outperformed the ball variant in most cases, which might be because boxes contain more parameters than balls when embedded in the same dimensional space.
Interestingly, on the MCG and Hearst datasets, region-based methods outperformed hyperbolic methods. This may be attributed to the relatively low “hyperbolicity” of these datasets, as indicated by their larger $\delta$-hyperbolicity values in Table 2. In contrast, on the Noun dataset, despite its high $\delta$-hyperbolicity value, hyperbolic methods performed better. This discrepancy is likely due to the larger dataset size—containing over twice the number of nodes compared to the others—which leads to crowding effects in Euclidean space that limit the effectiveness of region-based methods. Nevertheless, our method performed consistently well in both cases, as it can adjust the hyperbolic component by setting different $\lambda$ values in Equation (7).
Table 3: F1 score (%) for the inference task on ontologies.
Dataset Dim GALEN GO ANATOMY
ELBE d=5 20.7 36.9 43.1
    + $\tau$Box 20.8 32.2 42.2
    + RegD 25.2 50.0 58.7
ELBE d=10 21.2 42.4 43.0
    + $\tau$Box 20.7 34.7 47.2
    + RegD 25.8 50.5 62.5
ELEM d=5 16.9 23.5 34.6
    + RegD 19.2 36.4 52.5
ELEM d=10 17.3 27.4 38.7
    + RegD 18.8 40.0 55.5
6.2 Inference over Ontologies
Benchmark

### v2

@@ 3.1 Depth Dissimilarity: A Similarity Measure Incorporating Depth
It is important to recognize that evaluation metrics such as F1 scores and Hits@k rely solely on the ranking induced by the scoring function. As a result, only the relative order of the scores matters, while their absolute values are less important. Therefore, preserving the same ranking as that produced by hyperbolic distances is sufficient to attain comparable performance. Consequently, the depth-based dissimilarity defined in Equations 2 and 3, with the function $g(x)=x$, are well-suited for our implementation, as justified by Proposition 3.
3.2 Boundary Dissimilarity: A Non-Symmetric Measure of Inclusion
Although the depth dissimilarity $d_{\text{dep}}$ introduced above has been shown to have great power for embedding hierarchical data, it is a symmetric metric and therefore inadequate for fully capturing the inherently non-symmetric hierarchical relationships between objects. To address this limitation, we introduce the boundary dissimilarity, specifically designed to reflect the partial order of regions defined by set inclusion.
Our boundary dissimilarity generalizes the signed-distance-to-boundary in ShadowCone (Theorem 4.2, Yu et al. (2024)) by extending its applicability from specialized hyperbolic cone geometries to arbitrary Euclidean regions. This generalization is formally defined in Definition 2. As a result, our boundary dissimilarity can be computed in a much simpler form (Example 3) compared to the formulation used in ShadowCone (Equation 6).
Construction
The boundary dissimilarity is defined to measure the minimal cost associated with transforming the spatial relationship between two regions $\textit{reg}_{1}$ and $\textit{reg}_{2}$. Specifically, it quantifies the cost of moving $\textit{reg}_{2}$ out of $\textit{reg}_{1}$ when $\textit{reg}_{2}\subseteq\textit{reg}_{1}$, or moving $\textit{reg}_{2}$ into $\textit{reg}_{1}$ otherwise (when $\textit{reg}_{2}\not\subseteq\textit{reg}_{1}$). This cost can be defined for arbitrary geometric objects based on either distance or volume within Euclidean or other spaces. Below, we introduce a boundary dissimilarity based on the Euclidean distance for two regions $\textit{reg}_{1},\textit{reg}_{2}\subseteq\mathbb{R}^{n}$, which consists of two cases:
  1. 1.
    Containment ($\textit{reg}_{2}\subseteq\textit{reg}_{1}$): As illustrated in the left of Figure 3, when $\textit{reg}_{2}$ is fully contained within $\textit{reg}_{1}$, the boundary dissimilarity is defined by the minimum Euclidean distance between the complementary region $\textit{reg}_{1}^{c}$ and the points in $\textit{reg}_{2}$ (i.e., length of the red line). This distance quantifies the minimum translation cost required to move at least a part of $\textit{reg}_{2}$ out of $\textit{reg}_{1}$.
  2. 2.
    Non-Containment ($\textit{reg}_{2}\not\subseteq\textit{reg}_{1}$): As shown in the right of Figure 3, when $\textit{reg}_{2}$ is not fully contained within $\textit{reg}_{1}$, the boundary dissimilarity is defined as the maximum Euclidean distance from the points in $\textit{reg}_{2}\setminus\textit{reg}_{1}$ to $\textit{reg}_{1}$ (i.e., length of the red line). This distance quantifies the minimum translation cost for moving $\textit{reg}_{2}$ into $\textit{reg}_{1}$.
@@ 3.2 Boundary Dissimilarity: A Non-Symmetric Measure of Inclusion
Figure 2: Illustration of $d_{\text{bd}}(\textit{reg}_{1},\textit{reg}_{2})$ (red) for $\textit{reg}_{2}\subseteq\textit{reg}_{1}$ (left) or $\textit{reg}_{2}\not\subseteq\textit{reg}_{1}$ (right). Green lines shows the inverse: $d_{\text{bd}}(\textit{reg}_{2},\textit{reg}_{1})$.
$\textit{reg}_{2}$$\textit{reg}_{1}$$\textit{reg}_{1}$$\textit{reg}_{2}$$\textit{reg}_{3}$
[…]
Definition 2 (Boundary Dissimilarity).
Given a region space $\mathcal{R}$, we define the boundary dissimilarity over $\textit{reg}_{1},\textit{reg}_{2}\in\mathcal{R}$ by:
@@
Note that a negative sign is added to $d_{\text{bd}}(\textit{reg}_{1},\textit{reg}_{2})$ in the containment case ($\textit{reg}_{2}\subseteq\textit{reg}_{1}$) to clearly distinguish it from other cases. Moreover, the boundary dissimilarity is inherently asymmetric, that is, $d_{\text{bd}}(\textit{reg}_{1},\textit{reg}_{2})\neq d_{\text{bd}}(\textit{reg}_{2},\textit{reg}_{1})$ in general. For example, as illustrated in Figure 3, the boundary dissimilarity in the reverse order, $d_{\text{bd}}(\textit{reg}_{2},\textit{reg}_{1})$, corresponds to the length of the green dashed line, which differs from the red line representing $d_{\text{bd}}(\textit{reg}_{1},\textit{reg}_{2})$.
Moreover, for widely used geometric objects such as balls and boxes, the boundary dissimilarity can be computed efficiently using simple arithmetic operations.
@@
If $\textit{reg}_{1}=\textit{ball}(\mathbf{c}_{1},r_{1})$ and $\textit{reg}_{2}=\textit{ball}(\mathbf{c}_{2},r_{2})$, the boundary dissimilarity have the form:
@@
For the case of boxes, where $\textit{reg}_{1}=\textit{box}(\mathbf{c}_{1},\mathbf{o}_{1})$ and $\textit{reg}_{2}=\textit{box}(\mathbf{c}_{2},\mathbf{o}_{2})$, the boundary dissimilarity have the form:
[…]
@@
The following proposition demonstrates that our definition of the boundary dissimilarity effectively captures the inclusion relationship between two regions in two key aspects, as illustrated in Figure 3: (i) it identifies whether one region is (exactly) contained within another, and (ii) it enhances discrimination in inclusion chains, as smaller regions tend to have larger boundary dissimilaritys. This property is useful for distinguishing shallow children from deeper ones. It is worth noting that the proposition below applies to any regions, not only boxes or balls.
@@
For the boundary dissimilarity $d_{\text{bd}}$ in Definition 2, the following properties hold:
@@
    If $\textit{reg}_{3}\subseteq\textit{reg}_{2}\subseteq\textit{reg}_{1}$, then $d_{\text{bd}}(\textit{reg}_{1},\textit{reg}_{3})\leq d_{\text{bd}}(\textit{reg}_{1},\textit{reg}_{2}).$
Specific Constructions for Boxes or Balls
For specific geometric regions like balls and boxes, we can create specialized distance functions to measure the set-inclusion relationship based on their intrinsic geometric properties or established methods. Our framework accommodates these specialized metrics by allowing them to replace the general boundary dissimilarity function.
@@
    Volume-based dissimilarity for boxes: Since the volume of a box can be computed as the product of its offsets along different dimensions, we can define a partial distance based on volume:
@@
    Hyperbolic dissimilarity for balls: Yu et al. (2024) introduced a series of circular cones in hyperbolic space and defined a boundary dissimilarity based on the hyperbolic distance between the apex of these cones. By utilizing the natural mapping from balls to circular cones, we can derive a new boundary dissimilarity for balls as follows (see Appendix B for more details):
[…]
3.3 Training
For a given pair $(u,v)$, we define their energy as a weighted sum with weight $\lambda$ that balances the contributions of the hyperbolic-like depth dissimilarity:
@@ 3.3 Training
We say that $u$ is considered a parent of $v$ (i.e., $u\prec v$) if $E(u,v)\leq t$, where $t$ is a threshold that achieved the best performance on the evaluation set.
For model training, we use $d_{\text{dep}}$ from Equation 2 or 3 with the contrastive loss from Yu et al. (2024):
$\mathcal{L}(\gamma_{1},\gamma_{2})=\sum_{(u,v)\in P}\Bigg(\max\{E(u,v),\gamma_{1}\}+\log\bigg(\sum_{(u,v^{\prime})\in N}e^{\max\{\gamma_{2}-d_{\text{bd}}(\textit{reg}_{u},\textit{reg}_{v^{\prime}}),0\}}\bigg)\Bigg),$ (8)
where $P$ and $N$ denote positive and negative sample pairs, respectively. For positive pairs $(u,v)\in P$, the loss based on $E(u,v)$ promotes both the containment of $\textit{reg}_{v}$ within $\textit{reg}_{u}$ (via the $d_{\text{bd}}$ term) and their geometric similarity (via the $d_{\text{dep}}$ term), whose contributions are controlled by the weight $\lambda$ and the threshold $\gamma_{1}$. For negative pairs $(u,v^{\prime})\in N$, since our primary goal is to push $\textit{reg}_{v^{\prime}}$ outside of $\textit{reg}_{u}$, it is sufficient to use the boundary dissimilarity $d_{\text{bd}}(\textit{reg}_{u},\textit{reg}_{v^{\prime}})$ rather than the energy $E(u,v^{\prime})$. A threshold $\gamma_{2}$ is used to regulate how far $\textit{reg}_{v^{\prime}}$ is pushed from $\textit{reg}_{u}$.
@@
ShadowCone (Umbral-half) 57.7 69.4 45.2 52.2 36.8 40.1 32.8 32.6
(Penumbral-half) 52.8 67.8 44.6 51.7 35.0 37.6 26.8 28.4
RegD (box) 64.9 71.6 53.8 51.3 50.7 58.5 42.8 49.6
@@ 4.1 Inferences over DAG
Benchmark Following Yu et al. (2024), we evaluate our method on four real-world datasets consisting of Is-A relations: MCG Wang et al. (2015); Wu et al. (2012), Hearst patterns Hearst (1992), the WordNet Fellbaum (1998) noun taxonomy, and its mammal subgraph. All models are trained exclusively on basic edges, which are edges not implied transitively by other edges in the graph. For validation and testing, we use the same sets as in Yu et al. (2024), consisting of 5% of non-basic (inferred) edges, ensuring a fair comparison. The hyperparameter settings are provided in Appendix D.1. We exclude non-basic edges from training since they can be transitively derived from basic edges. Including them would artificially inflate performance metrics without properly evaluating the embeddings’ ability to capture hierarchical structures. For completeness, results for non-basic cases are provided in Appendix D.3.
@@
Baselines We compare our method RegD with (i) hyperbolic approaches such as EntailmentCone Ganea et al. (2018b) and ShadowCone Yu et al. (2024), which is the latest method with the state-of-the-art performance; (ii) region-based methods like OrderEmbedding Bordes et al. (2013), and tBox Boratko et al. (2021). We also compare with the ontology embedding methods, ELBE Peng et al. (2022) and ELEM Kulmanov et al. (2019), which can be considered as the baseline approaches embedding the DAG as boxes or balls, respectively. We use F1-scores as in previous studies.
@@
Results The performance comparison across different DAGs is shown in Table 2. RegD achieved the best performance on all four datasets. Notably, the box variant consistently outperformed the ball variant in most cases, which might be because boxes contain more parameters than balls when embedded in the same dimensional space. Interestingly, region-based methods outperformed hyperbolic methods on the MCG and Hearst datasets. However, on the Noun and mammal dataset, hyperbolic methods performed better. Nevertheless, our method performed consistently well in both cases, as it can adjust the hyperbolic component by setting different $\lambda$ values in Equation (7). The results of the ablation study are presented in Appendix D.2.
4.2 Inference and Link Prediction over Ontologies